# Unleashing the Power of LLMs as Multi-Modal Encoders for Text and Graph-Structured Data

### v1

@@ Abstract.
Graph-structured information offers rich contextual information that can enhance language models by providing structured relationships and hierarchies, leading to more expressive embeddings for various applications such as retrieval, question answering, and classification. However, existing methods for integrating graph and text embeddings, often based on Multi-layer Perceptrons (MLPs) or shallow transformers, are limited in their ability to fully exploit the heterogeneous nature of these modalities. To overcome this, we propose Janus, a simple yet effective framework that leverages Large Language Models (LLMs) to jointly encode text and graph data. Specifically, Janus employs an MLP adapter to project graph embeddings into the same space as text embeddings, allowing the LLM to process both modalities jointly. Unlike prior work, we also introduce contrastive learning to align the graph and text spaces more effectively, thereby improving the quality of learned joint embeddings. Empirical results across six datasets spanning three tasks—knowledge graph-contextualized question answering, graph-text pair classification, and retrieval—demonstrate that Janus consistently outperforms existing baselines, achieving significant improvements across multiple datasets, with gains of up to 11.4% in QA tasks. These results highlight Janus’s effectiveness in integrating graph and text data. Ablation studies further validate the effectiveness of our method.
Large Language Models, Representation Learning, Graph Neural Networks, Contrastive Learning
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
@@
In the realm of natural language processing (NLP), text embeddings play a pivotal role by transforming textual information into numerical representations, which facilitate a multitude of machine learning applications on the Web, such as question answering (QA) (Devlin et al., 2019; Reimers and Gurevych, 2019), retrieval (Ni et al., 2022a; Wang et al., 2022; Ma et al., 2024; Choi et al., 2021), and classification tasks (Wang et al., 2022; Devlin et al., 2019). These applications can benefit from the integration of graph-structured data to enhance the capabilities of NLP systems, by providing contextual information or by augmenting the original tasks with additional information. For example, prior work has shown that a QA system that includes a knowledge graph as input can leverage the relationships and hierarchies within the graph to more accurately understand and handle complex queries (Yasunaga et al., 2021; Zhang et al., 2022). To effectively integrate these two modalities, it is essential to develop methods to learn joint embeddings of graph-structured data and text data. Such embeddings can provide a unified representation that captures important information from both modalities, leading to performance improvements across various NLP tasks.
@@
Prior research has introduced various ways to learn joint embeddings of text and graph-structured data for embedding tasks (Yasunaga et al., 2021; Feng et al., 2020; Zhang et al., 2022; Lin et al., 2019a). These methods typically utilize either a Multi-layer Perceptron (MLP) or a shallow transformer (Vaswani et al., 2017) to integrate text features and graph embeddings encoded respectively by language models (LMs) (Devlin et al., 2019; Liu et al., 2019) and graph neural networks (GNNs) (Kipf and Welling, 2017; Velickovic et al., 2018; Xu et al., 2019). Despite their effectiveness, these approaches demonstrate a restricted capacity to fuse the features of the two modalities. The primary limitation arises from the limited ability of MLPs and shallow transformers to manage the high-dimensional and heterogeneous nature of joint embeddings, which can result in sub-optimal utilization of the rich contextual information from text and graph data. Recently, large language models (LLMs) have demonstrated significant potential for integrating and understanding modalities beyond just text. A representative example of this capability is in Vision-Language Models (VLMs) (Liu et al., 2023; Zhu et al., 2024a; Chen et al., 2023; Li et al., 2023), where visual tokens are combined with textual input and processed together by LLMs. This integration leverages the powerful capabilities of LLMs to handle multimodal data, allowing for a more holistic understanding of content that spans different forms of information. Inspired by this, we explore the potential of employing LLMs to better integrate text and graph-structured data, aiming to overcome the limitations observed from current approaches.
@@
In this paper, we present Janus, a simple yet effective framework to learn joint embeddings of text and graph data, leveraging the advanced capabilities of LLMs to address the limitations of previous approaches. Our method seamlessly integrates graph and text embeddings within the LLM framework, enhancing their alignment and interaction. Specifically, we transform graph embeddings into the same space as text embeddings using a multi-layer perceptron (MLP) adapter, enabling the LLM to process both modalities together. Additionally, we propose a contrastive learning strategy to better align the graph and text spaces, ensuring that the model learns richer representations of the combined data. Our extensive empirical analysis across a variety of NLP tasks highlights the key advantage of Janus: the ability to leverage LLMs’ strong language understanding and reasoning capabilities to process multimodal data, thus providing a more holistic and nuanced representation of both graph-structured and textual information. This integration leads to significant improvements in various NLP tasks, demonstrating the potential of Janus to push the boundaries of joint text and graph-embedding techniques.
@@
  • Integration of LLMs for Joint Embeddings: We propose Janus, a simple yet effective framework that leverages the strengths of LLMs to align and integrate graph and text embeddings. Janus effectively captures the rich contextual information of both modalities, enabling more robust joint representations.
  • Contrastive Learning for Graph-Text Alignment: We introduce a contrastive learning mechanism to explicitly align graph and text embeddings, enabling the model to better integrate the two modalities. This alignment enhances the quality of joint representations, leading to improved performance in downstream tasks.
  • Extensive Empirical Validation: We conduct extensive experiments on six datasets spanning three different tasks: knowledge graph (KG)-contextualized QA, graph-text pair classification, and retrieval tasks. Especially on the KG-contextualized QA tasks, we found that our Janus achieved an absolute accuracy improvement of 6.89% on CommonsenseQA, 9.46% on OpenBookQA, and 11.4% on MedQA. In graph-text pair classification and retrieval tasks, Janus also achieves superior performance by effectively integrating and aligning graph and text embeddings.
Figure 1. Overview of Janus framework. Unlike the common use of LLMs for generation tasks, we leverage LLMs to obtain joint embeddings of both text and graph data. We encode the input graph with a GNN, which provides the graph embeddings. The graph embeddings are then transformed into the word embedding space in the large language model. These embeddings are then fed into a large language model, and the outputs are utilized for various downstream tasks.
2. Problem Statement
Given an input text $x$ and its corresponding graph context $\mathcal{G}$, where $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$ consists of a set of nodes $\mathcal{V}$ and edges $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{R}\times\mathcal{V}$ that connect nodes via relationships $\mathcal{R}$, our goal is to extract joint embeddings $\phi(x,\mathcal{G})$. Here, $\phi$ is a learned function that maps the input text $x$ and graph $\mathcal{G}$ into a unified vector representation, capturing the multimodal information. These joint embeddings $\phi(x,\mathcal{G})$ are then used for downstream tasks. In this paper, we focus on three specific downstream tasks: multi-choice QA contextualized by knowledge graphs (KG), graph-text pair classification, and retrieval tasks.
2.1. KG-Contextualized QA
Given a question $q$, KG context $\mathcal{G}$, and answer candidate set with $n$ choices $\bm{a}=\{a_{1},\cdots,a_{n}\}$, KG-contextualized QA tasks aim to find the correct answer $a_{i}$ from $\bm{a}$. Each choice $a_{i}$ is first concatenated with the question $q$, leading to the input $x=[q,a_{i}]$, where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the concatenation of text or vectors. We then extract the joint embeddings $\phi(x,\mathcal{G})$ which is fed into a MLP layer to calculate scores. The choice with the highest score is selected as the prediction.
2.2. Graph-Text Pair Classification
@@
The significance of this task lies in its ability to improve the quality of training data for tasks such as generating text from knowledge graphs and vice versa (Ke et al., 2021; Lehmann et al., 2015; Colas et al., 2022). By accurately classifying graph-text pairs, the model can help reduce noise in training data, which in turn improves the overall performance of generation tasks (Lehmann et al., 2015).
2.3. Retrieval
For the retrieval task, given a query $q$ accompanied by its graph context $\mathcal{G}_{q}$, the goal is to retrieve the most relevant candidate from a set of options. Each candidate $c_{i}$ in the candidate set $\bm{c}=\{c_{1},\cdots,c_{m}\}$ also has an associated graph context $\mathcal{G}_{c_{i}}$. The task involves comparing the query-graph pair $(q,\mathcal{G}_{q})$ against each candidate-graph pair $(c_{i},\mathcal{G}_{c_{i}})$.
@@ 3. Methodology
In this work, we propose a simple yet effective framework Janus to learn joint embeddings of text and graphs, as illustrated in Figure 1. Specifically, graph-structured data are first extracted into a graph token, which is then fed into the LLM backbone together with the text tokens (§3.1, §3.2). The LLM backbone outputs the joint embeddings, which can be used for the downstream tasks, such as classification and retrieval.
@@
A key aspect of Janus is the explicit alignment between the graph and text embeddings. This alignment is crucial, as it allows the LLM backbone to better integrate the structured knowledge from the graph with the unstructured text, improving the quality of the joint representations (§4). To achieve this, we introduce a contrastive learning mechanism that explicitly maps embeddings from both modalities into a shared space (§3.3).
@@
While LLMs are commonly used for generation tasks in an auto-regressive paradigm, Janus takes a different route by utilizing LLMs as powerful encoders. This allows us to directly obtain robust joint embeddings of text and graph data by leveraging the rich contextual understanding of LLMs.
@@
Our proposed architecture, Janus, consists of three main components: a graph encoder, an MLP adapter, and a large language model backbone. Below, we discuss the components’ design and implementation details, along with the alignment mechanism.
3.1. Graph Data Encoding
@@
Janus employs a graph encoder to obtain graph embeddings, which encapsulate essential information extracted from the contextual structure of the graph. Next, we describe the two-step process, which includes: (1) the integration of the query node in the graph; and (2) the graph encoding.
3.1.1. Query Node Integration into Graph Structures
@@
We first initialize node embeddings within the input graph $\mathcal{G}$ with a language model (e.g., RoBERTa (Liu et al., 2019)). Following prior work (Lin et al., 2019b; Yasunaga et al., 2021; Zhang et al., 2022), we link entities mentioned in the query to nodes in the graph, denoting these nodes as $\mathcal{V}_{{\rm linked}}$. We then introduce a new node termed the query node, denoted as $v_{q}$. This query node is also initialized using a language model by encoding the input query text. The query node $v_{q}$ is then connected to all nodes within $\mathcal{V}_{{\rm linked}}$, enhancing the connection between the query and the nodes within the graph. We denote the updated graph with $\mathcal{G}^{\prime}=\{\mathcal{V}^{\prime},\mathcal{E}^{\prime}\}$.
3.1.2. Graph Encoding Process
@@
The updated graph $\mathcal{G}^{\prime}$ is then fed into an encoder for feature extraction. We adopt a modified version of graph attention network (GAT) (Velickovic et al., 2018; Yasunaga et al., 2021; Zhang et al., 2022) as the graph encoder. Specifically, in each layer of GAT, the message-passing process is formulated as
(1) $\bm{h}_{v}^{(\ell+1)}=f_{n}\left(\sum_{s\in\mathcal{N}_{v}\cup\{v\}}\alpha_{sv}\mathbf{m}_{sv}\right)+\bm{h}_{v}^{(\ell)}$
@@
(3) $\bm{q}_{s}=f_{q}(\bm{h}_{s}^{(\ell)},\bm{u}_{s})$ $\bm{k}_{v}=f_{k}(\bm{h}_{v}^{(\ell)},\bm{u}_{v},\bm{r}_{sv})$
(4) $\gamma_{sv}=\cfrac{\bm{q}_{s}^{\top}\bm{k}_{v}}{\sqrt{D}}$ $\alpha_{sv}=\cfrac{\exp(\gamma_{sv})}{\sum_{v^{\prime}\in\mathcal{N}_{s}\cup\{s\}}\exp(\gamma_{sv^{\prime}})}$
@@
3.2. Fusion of Graph and Textual Data in LLM
@@
Having obtained the graph embeddings, the next step involves integrating these with textual information using the LLM to produce embeddings suitable for downstream tasks. To facilitate this integration, we draw inspiration from practices in computer vision (Liu et al., 2023; Zhu et al., 2024a), where image embeddings are first transformed into the text space to be processed by LLMs. Similarly, we also convert the graph embeddings into the text space. We employ an MLP adapter for this purpose, which projects the graph token embeddings into the language model space using a two-layer MLP with ReLU activation functions, leading to the transformed graph embeddings $\tilde{\bm{g}}$. The transformation allows us to insert the processed graph token at the beginning of the text sequence, formatted as follows:
[…]
In this sequence, the graph token, output by the MLP adapter, precedes the textual tokens, which are derived from the initial input embeddings of the LLM. This arrangement ensures that the initial context for the LLM processing includes both graph-derived and textual information.
@@
We then feed this sequence into the LLM. The embeddings of the ¡$/s$¿ token from the last layer of the LLM are considered as the final output embeddings $\bm{z}$, encapsulating the combined knowledge of the graph and text inputs. This integration process leverages the LLM’s capacity to synthesize information across different modalities, optimizing the embeddings for subsequent applications.

### v2

# GT2Vec: Large Language Models as Multi-Modal Encoders for Text and Graph-Structured Data

@@ Abstract
Graph-structured information offers rich contextual information that can enhance language models by providing structured relationships and hierarchies, leading to more expressive embeddings for various applications such as retrieval, question answering, and classification. However, existing methods for integrating graph and text embeddings, often based on Multi-layer Perceptrons (MLPs) or shallow transformers, are limited in their ability to fully exploit the heterogeneous nature of these modalities. To overcome this, we propose GT2Vec, a simple yet effective framework that leverages Large Language Models (LLMs) to jointly encode text and graph data. Specifically, GT2Vec employs an MLP adapter to project graph embeddings into the same space as text embeddings, allowing the LLM to process both modalities jointly. Unlike prior work, we also introduce contrastive learning to align the graph and text spaces more effectively, thereby improving the quality of learned joint embeddings. Empirical results across six datasets spanning three tasks—knowledge graph-contextualized question answering, graph-text pair classification, and retrieval—demonstrate that GT2Vec consistently outperforms existing baselines, achieving significant improvements across multiple datasets. These results highlight GT2Vec’s effectiveness in integrating graph and text data. Ablation studies further validate the effectiveness of our method.
GT2Vec: Large Language Models as Multi-Modal Encoders for
Text and Graph-Structured Data
Jiacheng Lin1, Kun Qian2, Haoyu Han3, Nurendra Choudhary2, Tianxin Wei1, Zhongruo Wang2, Sahika Genc2, Edward W. Huang2, Sheng Wang4, Karthik Subbian2, Danai Koutra5, Jimeng Sun1 1University of Illinois Urbana-Champaign, 2Amazon, 3Michigan State University, 4University of Washington, 5University of Michigan
1 Introduction
@@
In the realm of natural language processing (NLP), text embeddings play a pivotal role by transforming textual information into numerical representations, which facilitate a multitude of machine learning applications on the Web, such as question answering (QA) Devlin et al. (2019); Reimers and Gurevych (2019), retrieval Ni et al. (2022b); Wang et al. (2022); Ma et al. (2024); Choi et al. (2021), and classification tasks Wang et al. (2022); Devlin et al. (2019). These applications can benefit from the integration of graph-structured data to enhance the capabilities of NLP systems, by providing contextual information or by augmenting the original tasks with additional information. For example, prior work has shown that a QA system that includes a knowledge graph as input can leverage the relationships and hierarchies within the graph to more accurately understand and handle complex queries Yasunaga et al. (2021); Zhang et al. (2022). To effectively integrate these two modalities, it is essential to develop methods to learn joint embeddings of graph-structured data and text data. Such embeddings can provide a unified representation that captures important information from both modalities, leading to performance improvements across various NLP tasks.
@@
Prior research has introduced various ways to learn joint embeddings of text and graph-structured data for embedding tasks Yasunaga et al. (2021); Feng et al. (2020); Zhang et al. (2022); Lin et al. (2019a). These methods typically utilize either a Multi-layer Perceptron (MLP) or a shallow transformer Vaswani et al. (2017) to integrate text features and graph embeddings encoded respectively by language models (LMs) Devlin et al. (2019); Liu et al. (2019) and graph neural networks (GNNs) Kipf and Welling (2017); Velickovic et al. (2018); Xu et al. (2019). Despite their effectiveness, these approaches demonstrate a restricted capacity to fuse the features of the two modalities. The primary limitation arises from the limited ability of MLPs and shallow transformers to manage the high-dimensional and heterogeneous nature of joint embeddings, which can result in sub-optimal utilization of the rich contextual information from text and graph data. Recently, large language models (LLMs) have demonstrated significant potential for integrating and understanding modalities beyond just text. A representative example of this capability is in Vision-Language Models (VLMs) Liu et al. (2023); Zhu et al. (2024a); Chen et al. (2023); Li et al. (2023), where visual tokens are combined with textual input and processed together by LLMs. This integration leverages the powerful capabilities of LLMs to handle multimodal data, allowing for a more holistic understanding of content that spans different forms of information. Inspired by this, we explore the potential of employing LLMs to better integrate text and graph-structured data, aiming to overcome the limitations observed from current approaches.
@@
In this paper, we present GT2Vec, a simple yet effective framework to learn joint embeddings of text and graph data, leveraging the advanced capabilities of LLMs to address the limitations of previous approaches. Our method seamlessly integrates graph and text embeddings within the LLM framework, enhancing their alignment and interaction. Specifically, we transform graph embeddings into the same space as text embeddings using a multi-layer perceptron (MLP) adapter, enabling the LLM to process both modalities together. Additionally, we propose a contrastive learning strategy to better align the graph and text spaces, ensuring that the model learns richer representations of the combined data. Our extensive empirical analysis across a variety of NLP tasks highlights the key advantage of GT2Vec: the ability to leverage LLMs’ strong language understanding and reasoning capabilities to process multimodal data, thus providing a more holistic and nuanced representation of both graph-structured and textual information. This integration leads to significant improvements in various NLP tasks, demonstrating the potential of GT2Vec to push the boundaries of joint text and graph-embedding techniques.
@@
  • Integration of LLMs for Joint Embeddings: We propose GT2Vec framework that leverages the strengths of LLMs to align and integrate graph and text embeddings. GT2Vec effectively captures the rich contextual information of both modalities, enabling more robust joint representations.
  • Contrastive Learning for Graph-Text Alignment: We introduce a contrastive learning mechanism to explicitly align graph and text embeddings, enabling the model to better integrate the two modalities.
  • Extensive Empirical Validation: We conduct extensive experiments on six datasets spanning three different tasks: knowledge graph (KG)-contextualized QA, graph-text pair classification, and retrieval tasks. GT2Vec achieves superior performance on all the three tasks, demonstrating its ability to effectively integrate graph and text data for enhanced multi-modal representation learning.
Figure 1: Overview of GT2Vec framework. Unlike the common use of LLMs for generation tasks, we leverage LLMs to obtain joint embeddings of both text and graph data. We encode the input graph with a GNN, which provides the graph embeddings. The graph embeddings are then transformed into the word embedding space in the large language model. These embeddings are then fed into a large language model, and the outputs are utilized for various downstream tasks.
2 Problem Statement
Given an input text $x$ and its corresponding graph context $\mathcal{G}$, where $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$ consists of a set of nodes $\mathcal{V}$ and edges $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{R}\times\mathcal{V}$ that connect nodes via relationships $\mathcal{R}$, our goal is to extract joint embeddings $\phi(x,\mathcal{G})$. Here, $\phi$ is a learned function that maps the input text $x$ and graph $\mathcal{G}$ into a unified vector representation, capturing the multimodal information. These joint embeddings $\phi(x,\mathcal{G})$ are then used for downstream tasks. In this paper, we focus on three specific downstream tasks: multi-choice QA contextualized by KG, graph-text pair classification, and retrieval tasks.
2.1 KG-Contextualized QA
Given a question $q$ in text form, KG context $\mathcal{G}$, and answer candidate set with $n$ choices $\bm{a}=\{a_{1},\cdots,a_{n}\}$, KG-contextualized QA tasks aim to find the correct textual answer $a_{i}$ from $\bm{a}$. Each choice $a_{i}$ is first concatenated with the question $q$, leading to the input $x=[q,a_{i}]$, where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the concatenation of text. We then extract the joint embeddings $\phi(x,\mathcal{G})$ which is fed into a MLP layer to calculate scores. The choice with the highest score is selected as the prediction.
2.2 Graph-Text Pair Classification
@@
The significance of this task lies in its ability to improve the quality of training data for tasks such as generating text from knowledge graphs and vice versa Ke et al. (2021); Lehmann et al. (2015); Colas et al. (2022). By accurately classifying graph-text pairs, the model can help reduce noise in training data, which in turn improves the overall performance of generation tasks Lehmann et al. (2015).
2.3 Retrieval
For the retrieval task, given a textual query $q$ accompanied by its graph context $\mathcal{G}_{q}$, the goal is to retrieve the most relevant candidate from a set of text-based options. Each candidate $c_{i}$ in the candidate set $\bm{c}=\{c_{1},\cdots,c_{m}\}$ also has an associated graph context $\mathcal{G}_{c_{i}}$. The task involves comparing the query-graph pair $(q,\mathcal{G}_{q})$ against each candidate-graph pair $(c_{i},\mathcal{G}_{c_{i}})$.
@@ 3 GT2Vec
In this work, we propose a simple yet effective framework GT2Vec to learn joint embeddings of text and graphs, as illustrated in Figure 1. Specifically, graph-structured data are first extracted into a graph token, which is then fed into the LLM backbone together with the text tokens (§3.1, §3.2). The LLM backbone outputs the joint embeddings, which can be used for the downstream tasks, such as classification and retrieval.
@@
A key aspect of GT2Vec is the explicit alignment between the graph and text embeddings. This alignment is crucial, as it allows the LLM backbone to better integrate the structured knowledge from the graph with the unstructured text, improving the quality of the joint representations (§4). To achieve this, we introduce a contrastive learning mechanism that explicitly maps embeddings from both modalities into a shared space (§3.3).
@@
While LLMs are commonly used for generation tasks in an auto-regressive paradigm, GT2Vec takes a different route by utilizing LLMs as powerful encoders. This allows us to directly obtain robust joint embeddings of text and graph data by leveraging the rich contextual understanding of LLMs.
@@
Our proposed architecture, GT2Vec, consists of three main components: a graph encoder, an MLP adapter, and a large language model backbone. Below, we discuss the components’ design and implementation details, along with the alignment mechanism.
3.1 Graph Data Encoding
@@
GT2Vec employs a graph encoder to obtain graph embeddings, which encapsulate essential information extracted from the contextual structure of the graph. Next, we describe the two-step process, which includes: (1) the integration of the query node in the graph; and (2) the graph encoding.
3.1.1 Query Node Integration into Graph Structures
@@
We first initialize node embeddings within the input graph $\mathcal{G}$ with a language model (e.g., RoBERTa Liu et al. (2019)). Following prior work Lin et al. (2019b); Yasunaga et al. (2021); Zhang et al. (2022), we link entities mentioned in the query to nodes in the graph, denoting these nodes as $\mathcal{V}_{{\rm linked}}$. We then introduce a new node termed the query node, denoted as $v_{q}$. This query node is initialized using a language model by encoding the input query text. The query node $v_{q}$ is then connected to all nodes within $\mathcal{V}_{{\rm linked}}$, enhancing the connection between the query and the nodes within the graph. We denote the updated graph with $\mathcal{G}^{\prime}=\{\mathcal{V}^{\prime},\mathcal{E}^{\prime}\}$.
3.1.2 Graph Encoding Process
@@
The updated graph $\mathcal{G}^{\prime}$ is then fed into an encoder for feature extraction. We adopt a modified version of graph attention network (GAT) Velickovic et al. (2018); Yasunaga et al. (2021); Zhang et al. (2022) as the graph encoder. Specifically, in each layer of GAT, the message-passing process is formulated as
$\bm{h}_{v}^{(\ell+1)}=f_{n}\left(\sum_{s\in\mathcal{N}_{v}\cup\{v\}}\alpha_{sv}\mathbf{m}_{sv}\right)+\bm{h}_{v}^{(\ell)}$
@@
3.2 Fusion of Graph and Textual Data in LLM
@@
We then integrate the graph embeddings with textual information using the LLM to produce embeddings suitable for downstream tasks. To facilitate this integration, we draw inspiration from practices in computer vision Liu et al. (2023); Zhu et al. (2024a), where image embeddings are first transformed into the text space to be processed by LLMs. Similarly, we also convert the graph embeddings into the text space. We employ an MLP adapter for this purpose, which projects the graph token embeddings into the language model space using a two-layer MLP with ReLU activation functions, leading to the transformed graph embeddings $\tilde{\bm{g}}$. The transformation allows us to insert the processed graph token at the beginning of the text sequence, formatted as [graph $[\text{{graph token}},\text{<$s$>},token\text{ 1},token\text{ 2},...,\text{<$/s$>}]$. In this sequence, the graph token, output by the MLP adapter, precedes the textual tokens, which are derived from the initial input embeddings of the LLM. This arrangement ensures that the initial context for the LLM processing includes both graph-derived and textual information.
@@
We then feed this sequence into the LLM. The embeddings of the <$/s$> token from the last layer of the LLM are considered as the final output embeddings $\bm{z}$, encapsulating the combined knowledge of the graph and text inputs. This integration process leverages the LLM’s capacity to synthesize information across different modalities, optimizing the embeddings for subsequent applications.
3.3 Graph-Text Alignment via Contrastive Learning